the westward and southward components of the velosity of a ship are 6.7 knots and 12.5 knots respectively. find the speed of the ship and the direction in which it is sailing.

Grade:

2 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Velocity in west direction:
v_{w} = 6.7 knots
Velocity in south direction:
v_{s} = 12.5 knots
Speed of ship:
S = \sqrt{v_{w}^{2}+v_{s}^{2}}
S = \sqrt{6.7^{2}+12.5^{2}}
S = \sqrt{201.14} = 14.182
Direction:
Moving at an angle w.r.t to west
tan^{-1}(\frac{125}{67})




Sher Mohammad IIT Delhi
9 years ago
for he westward and southward components of the velosity of a ship are 6.7 knots and 12.5 knots respectively.the speed of sheep is sqrt (6.72+ 12.52) and the direction is tan-1(6.7/12.5)